OpenStudy (anonymous):

I am still new to trigonometry.. Can I have clear explanation for this question please ? By considering the graph y= cos x, or otherwise, express the following in terms of cos x . (a) cos( 360-x) (b) cos (x+180)

OpenStudy (mokeira):

For cos(x+180) If you draw the graph and look at when x=0, then cos(x+180) = -1. Whereas cos(x) = 1. So cos(x+180) = - cos(x)

OpenStudy (math&ing001):

Have you seen the trigonometric circle yet ?

OpenStudy (anonymous):

No actually

OpenStudy (math&ing001):

This is what it looks like:  You'll notice that cos(x) and cos(360-x) have the same value.